# Forecasting intraday foreign exchange volatility with functional GARCH approaches

### v2

@@ Abstract
This paper seeks to analyse and predict conditional intraday volatility curves in FX markets using functional Generalised AutoRegressive Conditional Heteroscedasticity (GARCH) models. Remarkably, taking account of cross-dependency dynamics between the major currencies significantly improves intraday conditional volatility forecasting. Additionally, incorporating intraday bid-ask spread using a functional GARCH-X model further enhances predictability. The precise volatility forecasts motivate the construction of intraday Value-at-Risk (VaR). An intraday risk management application highlights that predicted intraday VaR curves can help mitigate dramatic losses in intraday trading strategies, showcasing their practical economic benefits.
@@ 1 Introduction
Modelling and forecasting foreign exchange (FX) market volatility has become increasingly important due to the growing FX trading volume and recent spikes in volatility, with implied volatility doubling (GBP), tripling (EUR), and quadrupling (JPY) for three reserve currencies previously renowned for their price stability in the first quarter of 2020.111BIS Triennial Central Bank Survey of Foreign Exchange and Over-The-Counter Derivatives Markets (Bank for International Settlements), www.bis.org/statistics/triennialrep/2022survey_guidelinesturnover.pdf; Chatham Financial Insights, www.chathamfinancial.com/insights/market-volatility-impacts-fx-markets) Elevated volatility in the FX market, such as this, at an inter-daily level can put corporate operations at risk by detrimentally affecting operating costs (Lessard and Lightstone, 1986). At an intraday level, it can expose intraday traders to risk by affecting trading profits. Forecasting FX volatility at the inter-daily level has received extensive attention in existing literature (see, for example, Lustig et al., 2011; Gabaix and Maggiori, 2015; Evans and Rime, 2019). In contrast, forecasting intraday volatility is rare. It presents additional challenges, particularly given our limited understanding of intraday volatility behaviour and the increasingly complex and risky movements in global FX markets. This paper aims to fill this gap by modelling and forecasting the intraday volatility of three of the world’s most active currency pairs, USD/EUR, USD/GBP, and USD/JPY, enhancing intraday risk management practices. This is important because a wide range of scholars emphasize the increasing demand for intraday risk management tools due to the move toward shorter-term holding patterns. Engle (2000) provides an econometric perspective on high-frequency data, highlighting the need for real-time risk management techniques to handle volatility and liquidity risks. Avellaneda and Lee (2010) discuss how high-frequency trading has increased the need for intraday risk management. Alexander and Kaeck (2008) address the market’s shift towards more dynamic, intraday models to manage sudden changes in risk, reflecting the evolving demands of practitioners in risk management.
@@
Forecasting volatility at an intraday level is difficult using traditional volatility models. The early work of Andersen and Bollerslev (1997) highlights the need to deploy new tools in the higher frequency environment by concluding that scalar GARCH-type models are not useful for modelling the dynamics of intraday observations. Subsequently, realised volatility-based models are developed to handle return volatility by incorporating intraday trading information (see, for example, Barndorff-Nielsen and Shephard, 2002; Sévi, 2014; Barunik et al., 2016). However, these approaches mainly forecast volatility on an inter-daily basis by capturing the dynamics of aggregated intraday volatility, making forecasting conditional volatility of returns at an intraday level difficult.
@@
To tackle the aforementioned challenge, this paper considers intraday return curves, deemed functional data objects, observed at an inter-daily frequency with the entire intraday movements well preserved. The conditional volatility of intraday return curves can be modelled and forecast using recently developed functional GARCH-type (FGARCH hereafter) models. The intuition for adopting the FGARCH-type models lies in their foundation in functional data analysis. For monographs on functional data analysis theory and applications, we refer interested readers to Bosq (2000), Ramsay and Silverman (2006), Ferraty and Vieu (2006), Horváth and Kokoszka (2012), Zhang (2014) and Kokoszka and Reimherr (2021). FGARCH-type models facilitate the modelling of volatility clustering dynamics in the FX market, incorporating the inherent second-moments of intraday trading movements (Cerovecki et al., 2019; Rice et al., 2020b). These models have been developed in the last decade to study the conditional heteroscedasticity of intraday return curves. Hörmann et al.’s (2013) seminal work proposes a functional autoregressive conditional heteroscedasticity (ARCH) model, which is subsequently generalised to the FGARCH(1,1) model in Aue et al. (2017) and the FGARCH$(p,q)$ model in Cerovecki et al. (2019). Rice et al. (2023) develops an FGARCH-X model to incorporate exogenous information with an additional X covariate. Using a goodness-of-fit test, Rice et al. (2020a) indicate that the FGARCH(1,1) model adequately explains the conditional heteroskedasticity of intraday return curves for many financial assets. Another recent study focusing on functional volatility includes Wang et al. (2023), who extend the functional volatility model of Müller et al. (2011) and highlight the dependence structure of intraday asset returns.
@@
To apply this model, we first investigate the dependence structure of intraday return curves in the FX market and assess whether the existing FGARCH-type models are suitable for leveraging the uncovered salient features for accurate forecasting. Specifically, we observe that the intraday return curves of USD/EUR, USD/GBP and USD/JPY exhibit strong cross-dependence, while in time series, they are uncorrelated and display long-range conditional heteroscedasticity. However, existing FGARCH-type models can only capture short-range dependence and explain the dynamics of the intraday volatility process in the univariate scenario, omitting their rich dependence structures, such as the long-range conditional heteroscedasticity and the cross-dependence structure between multiple assets.
@@
Our paper focuses on the aforementioned issues and tries to address them comprehensively. We develop a parsimonious FGARCH model incorporating covariance between assets using a multi-level functional principal component analysis (multi-level FPCA) during dimension reduction, which avoids extending heavily parameterized multivariate GARCH models into an FGARCH framework. Under the same model framework, we also introduce a long-range dependent functional principal component analysis (LFPCA) to account for the long-range conditional heteroscedasticity. Empirically, we are interested in overnight cumulative intraday return (OCIDR) curves. This return object is particularly interesting to study because (1) its smoothing feature makes it more suitable for functional time series modelling (Rice et al., 2020a), (2) it accounts for the overnight effect, which is crucial for volatility modelling (Hansen and Lunde, 2006b), and (3) it provides a convenient framework for modelling the volatility of inter-daily closing returns at each intraday grid point. To model the conditional volatility of OCIDR curves, we consider the FGARCH(1,1) model and also the functional GARCH-X (FGARCH-X) model (Rice et al., 2023) to incorporate intraday bid-ask spread microstructure information. Our findings show that taking account of the cross-dependencies between multiple currencies sharpens the predictability of the FGARCH model on the conditional volatility of intraday return curves. Additionally, incorporating intraday bid-ask spread information further improves FX volatility forecasting. With precise intraday volatility forecasts, we construct intraday value-at-risk (VaR) curves, offering risk management tools to mitigate sharp losses in a benchmark trading strategy during extreme market events.
@@
Our four main contributions are as follows. 1) We observe from the empirical data that the FX intraday returns are highly cross-dependent and exhibit long-range conditional heteroscedasticity. To accommodate these features in the existing FGARCH-type models, we propose two new data-driven basis methods for the dimension-reduction process in estimating the FGARCH-type models, namely multi-level FPCA and long-range FPCA methods, that preserve the properties of cross-dependence of the return curves and their long-range dependences, respectively. Both methods generally show improved out-of-sample performance compared with the standard static and dynamic FPCA dimensional reduction methods commonly used in the literature. This is particularly evident in the multi-level FPCA approach, which uses common and currency-specific factors in FX currencies. 2) We model and forecast the conditional volatility of FX intraday return curves through an FGARCH-X model. The FGARCH-X model, incorporating exogenous microstructure information – intraday bid-ask spreads, enhances volatility forecasting by improving explainability of long-range conditional heteroscedasticity and adjusting the level of volatility forecasts. Our empirical results suggest that, consistent with inter-daily volatility forecasting, the bid-ask spread information enhances the predictability of the intraday returns conditional volatility. 3) We statistically compare the relative performance of the volatility forecasting model candidates using the robust loss function approach of Patton (2011) combined with Diebold and Mariano’s (1995) test and the model confidence set (Hansen et al., 2011). This comparison was complicated by latent intraday volatility and the lack of a consistent estimator. 4) In order to explore the economic application of intraday volatility forecasting and to economically evaluate our out-of-sample forecasts, we compute intraday VaR using predicted values of conditional volatility. We uncover improvements in annual returns, Sharpe ratios, and maximum drawdowns using a VaR-informed intraday trading strategy. The trading strategy successfully avoids dramatic declines in intraday trading activities during extreme market conditions.
@@
The paper is structured as follows. We begin by highlighting the existing FGARCH-type models used to estimate and forecast the conditional volatility of intraday return curves in Section 2. Section 3 describes our dataset and discusses the properties of FX intraday returns and bid-ask spreads curves. Section 4 outlines the importance of dimension-reduction basis selection for adapting to FX intraday volatility forecasting. Section 5 presents the statistical forecasting performance of the models, and Section 6 provides intraday risk management applications. We conclude and make a recommendation for future work in Section 7.
2 Functional GARCH-type models
[…]

### v3

# Intraday FX Volatility-Curve Forecasting with Functional GARCH Approaches

@@ Abstract
This paper seeks to forecast intraday volatility curves for major foreign exchange (FX) currencies using functional GARCH models. Intraday return curves are observed at a daily frequency, yet preserve the full high-frequency trading structure, enabling volatility analysis at the intraday level. We demonstrate that the USD/EUR, USD/GBP, and USD/JPY intraday return curves exhibit strong cross-dependence, while individually they are serially uncorrelated but display long-range conditional heteroskedasticity. Embedding cross-currency dependence via multi-level functional principal component analysis and adding intraday bid-ask spread curves as exogenous drivers significantly improves intraday and day-ahead volatility forecasts relative to functional and realised-volatility baselines. The precise volatility forecasts motivate the construction of intraday Value-at-Risk (VaR). An intraday risk management application highlights that predicted intraday VaR curves can help mitigate dramatic losses in intraday trading strategies, showcasing their practical economic benefits in FX markets.
@@
Highlights
  • Forecast intraday FX volatility curves via functional GARCH-type models
  • Intraday FX curves display strong cross-dependence and long-range conditional heteroskedasticity
  • Modelling cross-currency dependence significantly improves forecast accuracy
  • Intraday bid-ask spread curves add predictive power
  • Intraday VaR curves provide a simple intraday risk management tool
@@ 1 Introduction
The modelling and forecasting of foreign exchange (FX) market volatility has become increasingly important due to the growing volume of FX trading and recent volatility spikes, with implied volatility doubling (GBP), tripling (EUR), and quadrupling (JPY) for three reserve currencies previously renowned for their price stability in the post-COVID period. [^1] BIS Triennial Central Bank Survey of Foreign Exchange and Over-The-Counter Derivatives Markets (Bank for International Settlements), https://www.bis.org/publ/qtrpdf/r_qt2212f.pdf; Chatham Financial Insights, www.chathamfinancial.com/insights/market-volatility-impacts-fx-markets). Downside risks and volatility spillovers across currencies present significant challenges for currency trading, particularly as activity shifts toward shorter intraday horizons (Chuliá et al.,, 2018; Rubaszek et al.,, 2025). Early studies document the growing demand for intraday risk management tools for high-frequency trading (see, e.g., Engle,, 2000; Alexander and Kaeck,, 2008; Avellaneda and Lee,, 2010). However, for decades, this demand remained largely unmet, as the field lacked suitable econometric tools and the literature primarily focused on inter-daily risk management. The primary challenge lies in forecasting FX volatility on the intraday horizon. This paper aims to enhance intraday risk management practices by proposing a methodological framework for forecasting the intraday volatility of three major currency pairs: USD/EUR, USD/GBP, and USD/JPY.
@@
While forecasting FX volatility at the inter-daily horizon has been extensively studied (see, e.g., Martens,, 2001; Gabaix and Maggiori,, 2015; Barunik et al.,, 2016), research on intraday volatility forecasting remains limited, as traditional models struggle to capture dynamics at such high frequencies. Early work by Andersen and Bollerslev, (1997) highlights the need for new approaches in this setting, concluding that standard GARCH-type models are inadequate for modelling intraday dynamics. Subsequently, realised volatility-based models have been developed to incorporate intraday trading information (see, for example, Barndorff-Nielsen and Shephard,, 2002; Sévi,, 2014; Barunik et al.,, 2016). Nonetheless, these models still primarily forecast volatility at the inter-daily level by aggregating intraday information, rather than modelling intraday volatility directly.
@@
To tackle this challenge, we consider FX intraday return curves, deemed functional data objects, observed at an inter-daily frequency with the entire intraday movements well preserved. The conditional volatility of intraday return curves can be modelled and forecast using recently developed functional GARCH-type (FGARCH hereafter) models. Intuitively, the conditional volatility inherits the functional form of the return curves, such that volatility itself is represented as an intraday curve, preserving fluctuations across the entire intraday trading horizon. The adoption of FGARCH-type models is motivated by their strong foundations in functional data analysis. [^2] For monographs on functional data analysis theory and applications, we refer interested readers to Bosq, (2000), Ramsay and Silverman, (2006), Horváth and Kokoszka, (2012), and Kokoszka and Reimherr, (2021). Variants of FGARCH-type models have been developed over the past decade to capture volatility clustering dynamics in financial markets (Hörmann et al.,, 2013; Aue et al.,, 2017; Cerovecki et al.,, 2019; Rice et al., 2020b, ; Rice et al.,, 2023). [^3] Hörmann et al.,’s (2013) seminal work proposes a functional autoregressive conditional heteroscedasticity (ARCH) model, which is subsequently generalised to the FGARCH(1,1) model in Aue et al., (2017) and the FGARCH(p,q)(p,q) model in Cerovecki et al., (2019). Rice et al., (2023) develops an FGARCH-X model to incorporate exogenous information with an additional X covariate.
@@
As intraday dynamics often exhibit complex comovements across global markets, direct applications of FGARCH models to the FX market may overlook important characteristics of FX volatility (Gabaix and Maggiori,, 2015; Rubaszek et al.,, 2025). We find that the intraday return curves of USD/EUR, USD/GBP, and USD/JPY, exhibit strong cross dependence; however, when viewed as isolated time series, they are serially uncorrelated and display long-range conditional heteroskedasticity. Existing FGARCH-type models can only capture short-range dependence and explain the dynamics of the intraday volatility process in the univariate scenario, omitting their rich dependence structures, such as the long-range conditional heteroskedasticity and cross-asset dependence in FX markets.
Our paper attempts to address the aforementioned methodological issues comprehensively. We develop a parsimonious FGARCH model that incorporates cross-currency connectedness by employing multi-level functional principal component analysis (MFPCA) for dimension reduction. This approach avoids the need to extend heavily parameterized multivariate GARCH models into the functional time series framework. Instead, we retain the existing FGARCH-type specifications but estimate the models through a dimension-reduction procedure based on basis functions derived from MFPCA across multiple FX assets, thereby incorporating cross-dependence. Under the same estimation framework, we also introduce a long-range dependent functional principal component analysis (LFPCA) to account for the long-range conditional heteroskedasticity in FX assets. Both methods demonstrate superior performance in forecasting FX volatility.
@@
Empirically, we forecast the intraday volatility of return curves, specifically the overnight cumulative intraday return (OCIDR) curves for three major FX pairs: USD/EUR, USD/GBP, and USD/JPY. This return object is particularly interesting to study because (1) its smoothing feature makes it more suitable for functional time series modelling (Rice et al., 2020a, ), (2) it accounts for the overnight effect, which is crucial for volatility modelling (Hansen and Lunde, 2006b, ), and (3) it provides a convenient framework for modelling the volatility of inter-daily closing returns at each intraday grid point. To model the conditional volatility of the OCIDR curves, we consider the FGARCH(1,1) model and the functional GARCH-X (FGARCH-X) model (Rice et al.,, 2023), which incorporates intraday bid-ask spread microstructure information. Our findings show that taking into account the connectedness between multiple currencies sharpens the predictability of the conditional volatility of intraday return curves. Additionally, incorporating intraday bid-ask spread information further improves FX volatility forecasting.
Building on precise intraday volatility forecasts, we introduce an intraday risk management framework by constructing intraday value-at-risk (VaR) curves. This approach mitigates sharp losses by reducing the maximum drawdown of a benchmark intraday trading strategy during extreme market events. The gains are strongest in high-uncertainty states, consistent with recent evidence on intraday connectedness and bad-shock asymmetries in FX markets.
@@
Our four main contributions are as follows: 1) We observe from the empirical data that the FX intraday returns are highly cross-dependent and exhibit long-range conditional heteroskedasticity. To accommodate these features in the existing FGARCH-type models, we propose two new data-driven basis methods for the dimension-reduction process in estimating the FGARCH-type models, namely, multi-level FPCA and long-range FPCA methods, which preserve the properties of cross-dependence of the return curves and their long-range dependence, respectively. Both methods generally show improved out-of-sample performance compared to the standard static and dynamic FPCA dimensional reduction methods commonly used in the literature. This is particularly evident in the multi-level FPCA approach, which uses common and currency-specific factors in FX currencies. 2) We model and forecast the conditional volatility of the intraday FX return curves through an FGARCH-X model. The FGARCH-X model, which incorporates exogenous microstructure information – intraday bid-ask spreads, enhances volatility forecasting by improving the explainability of long-range conditional heteroskedasticity and adjusting the level of volatility forecasts. Our empirical results suggest that, consistent with inter-daily volatility forecasting, the bid-ask spread information enhances the predictability of the intraday returns’ conditional volatility. 3) We statistically compare the relative performance of the volatility forecasting model candidates using the robust loss function approach of Patton, (2011) combined with Diebold and Mariano,’s (1995) test and the model confidence set of Hansen et al., (2011). This comparison was complicated by latent intraday volatility and the lack of a consistent estimator. 4) To explore the risk management implications of intraday volatility forecasting and to economically evaluate our out-of-sample forecasts, we compute intraday VaR using predicted values of conditional volatility. We uncover improvements in annual returns, Sharpe ratios, and maximum drawdowns using a VaR-informed intraday trading strategy. The trading strategy successfully avoids dramatic declines in intraday trading activities during extreme market conditions.
@@
The paper is structured as follows. We begin by describing our dataset and discussing the properties of FX intraday returns and bid-ask spreads curves in Section 2. Section 3 outlines existing FGARCH-type models and introduces new data-driven bases in dimension reduction for adapting FX intraday volatility forecasting. Section 4 presents the statistical forecast performance of the models, and Section 5 provides intraday risk management applications. We conclude and make a recommendation for future work in Section 6.
2 FX data and preliminary analysis
This section introduces the dataset and presents preliminary analyses of its properties. We consider the three most traded currencies against USD, namely EUR, GBP, and JPY. These currencies constituted over 75% of the turnover of OTC FX instruments in 2019, according to the Bank for International Settlements. [^4] Bank of International Settlements, https://stats.bis.org/statx/srs/table/d11.3 For each currency, the intraday spot, bid, and ask rates are collected at a 5-minute frequency from a collection of databases, including Morningstar, FXCM, TenFore, and Thomson Reuters. The sample ranges from 07-January-2014 to 30-September-2020.555Due to limited availability of high-frequency bid-ask spread data, the analysis is not extended to more recent periods, where FGARCH-X models with more recent bid-ask spread covariates would be required. In unreported results using a more recent sample, FGARCH models that only require intraday returns yield findings consistent with those reported here. The chosen sample, therefore, aligns to ensure comparability between the FGARCH and FGARCH-X specifications. Given the OTC nature of FX markets, our sample covers U.S. trading hours from midnight on Monday until the close on Thursday night. Intraday data on Sunday and Friday are truncated from our sample, as they do not form 24-hour trading sessions. Consequently, each FX rate forms an information set of 1,275 trading days, each with 288 intraday grid points. In an unreported analysis, we also include forecasting during the U.S. main trading session from 7:00 to 17:00, Monday to Friday, spanning 1,722 trading days with 108 daily grid points each day. The results show consistency with the dataset used in this study.
$y_{t}(u)=[\ln P_{t}(u)-\ln P_{t-1}(1)]\times 100,$ (2.1)
where $P_{t}(u)$ denotes the spot rate of either currency on day $t$ evaluated at intraday time $u$, $u\in[0,1]$, and $P_{t-1}(1)$ is the closing spot rate on day t1t-1. The functional data objects can be obtained using smoothing techniques (Ramsay and Silverman,, 2006, Chapter 3) or simple linear interpolation under the condition of continuity (Rice et al., 2020a, ). The OCIDR curves provide a continuous trajectory, so they are suitable for casting as functional time series objects while accounting for the overnight effect (Hansen and Lunde, 2006b, ). Desirably, forecasting the volatility of OCIDR curves includes inter-daily volatility prediction because OCIDR curves end at daily closing returns, defined as $y_{t}=\ln P_{t}(1)-\ln P_{t-1}(1)$.
Additionally, we obtain intraday bid-ask spread curves. The inclusion of the bid-ask spread is motivated by Evans and Rime, (2019), who point out that although the FX rate is commonly represented as the midpoint between the best bid and the ask price, in reality, all FX trades are executed at the bid or the ask price. This bid or ask trading dynamic also brings about imperfect measurement in determining the variance of return, further motivating our functional characterisation of the intraday data. The use of bid-ask spreads has been widely discussed in the volatility forecasting literature. For example, Bollerslev and Melvin, (1994) state that bid-ask spreads act as a measure of liquidity and have a significant relationship with the volatility of the FX rate. Incorporating bid-ask spread data also helps address the issue of bid-ask bounce, which might be misinterpreted as volatility arising from price movements. As a measure of liquidity, we calculate the intraday bid-ask spread curves (IBAS) through
$\text{IBAS}_{t}(u)=\text{ask}_{t}(u)-\text{bid}_{t}(u),\quad u\in[0,1].$
To achieve stationarity and generate an information flow that is compatible with the OCIDR curve, we calculate the overnight cumulative IBAS (OCIBAS) curve using $\text{OCIBAS}_{t}(u)=\text{IBAS}_{t}(u)-\text{IBAS}_{t-1}(1)$. The OCIBAS curve is also assumed to be square integrable with the sample path in $\mathcal{L}^{2}[0,1]$. Intuitively, higher values of the $\text{OCIBAS}_{t}(u)$ curve represent lower liquidity and vice versa.
Figure 2.1 displays the plots of the OCIDR and OCIBAS curves for USD/EUR. Each OCIDR curve records the evolution of the return movements over the trading day. The plot of the OCIBAS curves shows that they spike during opening and closing trading hours, implying less liquidity during these trading sessions. To understand the properties of these curve processes, we deploy the KPSS-type stationary test (Horváth et al.,, 2014), the static FPCA-type normality test (Górecki et al.,, 2018), the independence test (Kokoszka et al.,, 2017), and the conditional heteroskedasticity test (Rice et al., 2020b, ). Table 2.1 documents the p-values for each statistical test in the first training sample, which runs from 03-January-2014 to 31-December-2016. We limit our testing to the first training sample for in-sample analysis throughout the paper, and to ensure the robustness of our reported results, we usually replicate the analysis for other training samples. Table 2.1 indicates that the OCIDR curves are weakly stationary, non-Gaussian, serially uncorrelated, and conditionally heteroscedastic functional time series. Interestingly, the properties of the OCIDR curves uncovered here for the FX market are consistent with intraday return curves in the equity and commodity markets (see, e.g., Rice et al.,, 2023).
Table 2.1: $P$-values of statistical tests on Overnight Cumulative Intraday Return (OCIDR) curves and Overight Cumulative Intraday Bid-Ask Spread (OCIBAS) curves, as defined in Section 2, for our three currency pairs, USD/EUR, USD/GBP, and USD/JPY. The results are calculated for 07-January-2014 to 17-May-2017, the first training sample period of our data. We test for stationarity, normality, autocorrelation, and heteroskedasticity with the $\widehat{a}_{x}$ and $\widehat{a}_{x^{2}}$$\widehat{a}_{|x|}$ representing the short- and long-range dependency memory parameter estimates obtained using the Local Whittle (LW) estimator. Lags of 1, 5, 10, and 20 days are specified for the autocorrelation and heteroskedasticity tests.
Stationarity Normality Autocorrelation Heteroskedasticity $\widehat{a}_{x}$ $\widehat{a}_{x^{2}}$ Lag 1 5 10 20 1 5 10 20 $\mbox{OCIDR}_{\text{USD/EUR}}$ 0.22 0.00 0.71 0.61 0.43 0.60 0.04 0.00 0.00 0.00 0.02 0.20 $\mbox{OCIDR}_{\text{USD/GBP}}$ 0.60 0.00 0.64 0.68 0.56 0.40 0.06 0.00 0.00 0.00 -0.04 0.33 $\mbox{OCIDR}_{\text{USD/JPY}}$ 0.68 0.00 0.19 0.10 0.28 0.25 0.00 0.02 0.04 0.02 0.06 0.13 $\widehat{a}_{|x|}$ $\mbox{OCIBAS}_{\text{USD/EUR}}$ 0.11 0.00 0.37 0.25 0.36 0.37 0.96 0.09 0.01 0.00 0.11 0.12 $\mbox{OCIBAS}_{\text{USD/GBP}}$ 0.04 0.00 0.14 0.37 0.26 0.20 0.95 0.07 0.00 0.00 0.18 0.19 $\mbox{OCIBAS}_{\text{USD/JPY}}$ 0.07 0.00 0.61 0.83 0.34 0.34 0.00 0.00 0.00 0.00 0.10 0.27
We also investigate the long-range dependence structure of the level and squares of the FX curve sequences. According to Li et al., (2020, 2021), the memory parameter $a$ of the stationary curves can be estimated using either via a semiparametric R/S method or a feasible local Whittle (LW) estimator. Taking into account the efficiency of the estimator, we apply the latter method as suggested by Li et al., (2021), with the LW estimator $\widehat{a}_{x}$ denoting the memory parameter of the level curves, and $\widehat{a}_{x^{2}}$ denoting the memory parameter of the squared process. Estimates of $\widehat{a}_{x}$ and $\widehat{a}_{x^{2}}$ show that the levels of the OCIDR curves are short-range dependent, while the squared processes are long-range dependent. The absolute processes $\widehat{a}_{|x|}$ of OCIBAS curves are also long-range dependent.
Furthermore, it has been documented that co-movements between excess returns of a variety of internationally traded financial assets are widespread (see, e.g. Bekaert,, 1995). These studies motivate a series of works on forecasting daily volatility in the FX market. Recently, theoretical bases for understanding cross-currency co-movements have been built by Gabaix and Maggiori, (2015) and Colacito et al., (2018), with Gabaix and Maggiori, (2015) further highlighting the role of co-movements in modelling FX volatility. This cross-dependence structure between related currency pairs may also add predictability to return volatility in FX markets. Therefore, we seek to understand the interdependency between currency pairs by estimating the sample covariance operator between their OCIDR curves.
Figure 2.2 displays the estimated covariance operator between the various pairwise OCIDR curves. The results show that the three pairs of OCIDR curves are correlated, with a particularly high correlation between USD/EUR and USD/GBP, consistent with the dynamic connectedness of Rubaszek et al., (2025). The source of this cross-dependency can be attributed to shocks to the U.S. economy and U.S. interest rates, which, given the base currency of USD, impact all FX rates we consider. Eichenbaum and Evans, (1995) provide empirical evidence that U.S. monetary policy shocks, such as interest rate changes, have significant spillover effects on exchange rates globally, highlighting how U.S. economic shifts create dependencies among various currencies. Hausman and Wongswan, (2011) examine the impact of U.S. Federal Reserve announcements, particularly those concerning interest rates, on global asset prices, including exchange rates. This study supports the idea that cross-dependencies in currency markets often arise from U.S. economic shocks. Thus, the empirical finding of strong covariance motivates our position that accounting for cross dependence may enhance volatility modelling and forecasting.
3 Methodology
3.1 Functional GARCH-type models
This section outlines the existing functional GARCH-type models that are used to fit and forecast the conditional heteroskedasticity of intraday return curves, denoted as $\{y_{t}(u)\}$, where $u\in[0,1]$ represents the intraday time index, and $1\leq t\leq N$ denotes the inter-daily time index. We assume that the intraday return curve, $y_{t}(u)$, is square-integrable and constitutes sample paths in a $\mathcal{L}^{2}[0,1]$ Hilbert space. In addition, it is equipped with an inner product $\left\langle y_{1},y_{2}\right\rangle=\int y_{1}(u)y_{2}(u)du$, resulting in the norm $y(u)||=[\int y^{2}(u)du]^{1/2}$, for $\int\equiv\int_{0}^{1}$. For convenience, we also denote the integral operator of the kernel $\mathbf{g}(y)(u)=\int g(u,v)y(v)dv$, for $g(u,v)\in\mathcal{L}^{2}[0,1]^{2}$, and further define subspaces $\mathcal{L}^{2}[0,1]^{+}=\{y\in\mathcal{L}^{2}[0,1],y(u)\geq 0\}$ and $\mathcal{L}^{2}[0,1]_{*}^{+}=\{y\in\mathcal{L}^{2}[0,1],y(u)>0\}$, for almost every $u\in[0,1]$.